OpenStudy (anonymous):

What is √32/4 expressed in simplest radical form? a) √2 b) 4√2 c) √8 d) √8/2

OpenStudy (anonymous):

it's c

OpenStudy (anonymous):

\[\sqrt{2}\]

OpenStudy (anonymous):

or even simpler [2\sqrt{2}\]

OpenStudy (anonymous):

or even simpler \[2\sqrt{2}\]

OpenStudy (anonymous):

\[\frac{\sqrt{16*2}}{4}\Rightarrow \frac{\sqrt{16}*\sqrt{2}}{4}\Rightarrow \frac{4\sqrt{2}}{4}\Rightarrow \sqrt{2}\]

OpenStudy (anonymous):

ok, I misunderstood the problem

OpenStudy (anonymous):

a)